$SPAD/src/input c06aef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
x:Matrix SF:=
 [[0.34907 ,0.54890 ,0.74776 ,0.94459 ,1.13850 ,1.32850 ,1.51370 ]]
n:=ncols x
result:=c06eaf(n,x,-1)

———-

References^{[1] nothing}

4